# Online Feedback Optimization over Networks: A Distributed Model-free Approach

### v1

@@ I Introduction
The above OFO controllers are mostly centralized. When applied to large-scale networked systems involving interconnected subsystems, they can face issues of scalability, robustness, privacy disclosure, that are common in large-scale centralized decision architectures [15]. First, a central processor is needed to communicate with all the subsystems and generate network-wide control inputs. Thus, there exist severe computational and communication bottlenecks for sizable networks. Second, this centralized architecture is not robust against potential failures or attacks at the central unit. Finally, since gradient-based OFO controllers (e.g., [10, 11, 1, 7]) require the system-wide input-output sensitivity to perform local iterative updates, concerns on privacy disclosure and challenges of local modeling also arise.
@@ II-A Problem Formulation
We consider a fast stable networked system described by its steady-state nonlinear algebraic map $h:\mathbb{R}^{N}\times\mathbb{R}^{N}\rightarrow\mathbb{R}^{N}$
@@
where $u\in\mathbb{R}^{N}$ is the input, $y\in\mathbb{R}^{N}$ is the output, and $d\in\mathbb{R}^{N}$ is an unknown exogenous disturbance. The $i$-th elements of $u$, $y$, and $d$ are the local input, output, and disturbance of agent $i$, respectively, where $i=1,\ldots,N$. In general, the map (1) represents the steady-state input-output map of a nonlinear system with coupled dynamics between agents as follows
$\begin{split}x_{k+1}&=f(x_{k},u,d),\\ y_{k}&=g(x_{k},u,d),\end{split}$ (2)
where $x_{k}\in\mathbb{R}^{N}$ is the state, and $f$ and $g$ are functions in the dynamics equations. We focus on a per-agent scalar setting to simplify notation, however, our results readily extend to scenarios involving multiple variables.
@@
In problem (3), the global objective function $\Phi:\mathbb{R}^{N}\times\mathbb{R}^{N}\rightarrow\mathbb{R}$ is the average of all the local objectives $\Phi_{i}:\mathbb{R}\times\mathbb{R}\rightarrow\mathbb{R}$, and $u(i)$ and $y(i)$ are the input and output of agent $i$, respectively. Moreover, $\mathcal{U}_{i}\subseteq\mathbb{R}$ is the local input constraint set of agent $i$. The overall constraint set $\mathcal{U}$ is the Cartesian product of all the local sets. By incorporating (3b) into the objective $\Phi$, problem (3) can be transformed to:
@@
where $\tilde{\Phi}(u)=\frac{1}{N}\sum_{i=1}^{N}\tilde{\Phi}_{i}(u)$, and $\tilde{\Phi}_{i}(u)\triangleq\Phi_{i}(u(i),y(i))$ is the reduced local objective. The disturbance $d$ becomes an unknown parameter in $\tilde{\Phi}(u)$. We make the following assumptions on the objectives and the network.
@@
Both Assumptions 1 and 2 are standard in the literature of network control, see [28, 29, 27] and [15, 30], respectively.
@@ III Design of Distributed Model-free OFO Controllers
In contrast, we pursue an OFO controller to regulate the system (1) in a distributed and collaborative manner. This controller employs a novel communication protocol, where each agent stores and exchanges a vector of historical evaluations of local objectives to track the global objective value. This vector is updated via average consensus in the same time scale as the local optimization iterations. In the remainder of this paper, we treat the unconstrained case of (4) (i.e., $\mathcal{U}=\mathbb{R}^{N}$) separately, to derive sharper results.
@@ III-A Unconstrained Setting
4:        $z_{i,k}(l)=\sum_{j\in\mathcal{N}_{i}}W_{ij}\bm{z}_{j,k}(l)$, $\forall l=1,\ldots,\tau$;
5:        $z^{(1)}_{i,k}=[z^{\top}_{i,k},\tilde{\Phi}_{i}(u_{k}+\delta v_{k})]^{\top}$;
@@
$\Delta_{k}(i)\!=\!\begin{cases}\bm{z}^{(1)}_{i,k}(1)v^{\text{init}}_{0}(i)&\text{if $k\!=\!0$},\\ (\bm{z}^{(1)}_{i,k}(1)\!-\!\bm{z}^{(1)}_{i,k-1}(1))v^{\text{init}}_{k}(i)&\text{if $0\!<\!k\!<\!\tau$},\\ (\bm{z}^{(1)}_{i,k}(1)\!-\!\bm{z}^{(1)}_{i,k-1}(1))v_{k-\tau}(i)&\text{if $k\!\geq\!\tau$};\end{cases}$ (6)
7:        $z_{i,k+1}=[z^{(1)}_{i,k}(2),\ldots,z^{(1)}_{i,k}(\tau+1)]^{\top}$;
@@
The model-free distributed controller requires a suitable initialization of $\bm{z}_{i,0}\in\mathbb{R}^{\tau}$, where $\tau\in\mathbb{N}$ is a design parameter. Agent $i$ can evaluate the local objective $\tau$ times using $v^{\text{init}}_{l}(i)\sim\mathcal{N}(0,1)$ for $l=0,\ldots,\tau-1$ and stack the result in $\bm{z}_{i,0}$. Consequently, we have
@@
We summarize the detailed steps of the proposed distributed OFO controller in Algorithm 1. At every iteration $k$, each agent $i$ generates a random exploration signal $v_{k}(i)\sim\mathcal{N}(0,1)$ synchronously and evaluates its local objective value $\tilde{\Phi}_{i}(u_{k}+\delta v_{k})=\Phi_{i}(u_{k}(i)+\delta v_{k}(i),y_{k}(i))$ with its local input $u_{k}(i)+\delta v_{k}(i)$ and local measured output $y_{k}(i)$, which captures the collective effect of the coupled dynamics. These local objective values are stored in a vector $\bm{z}_{i,k}\in\mathbb{R}^{\tau}$ of size $\tau$. Then, the agents update $\bm{z}_{i,k}$ via the average consensus step (see line 4 in Algorithm 1), where $W$ is a doubly stochastic matrix with Metropolis weights[15]. Afterward, the new local evaluation of $\tilde{\Phi}_{i}$ queues into $\bm{z}^{(1)}_{i,k}$, and the first element of $\bm{z}^{(1)}_{i,k}$ is taken out to facilitate the optimization iterations (see lines 6-7 in Algorithm 1). We illustrate the process of communicating and updating $\bm{z}_{i,k}$ in Fig. 1.
In fact, when $k\geq\tau+1$, the first element of $\bm{z}^{(1)}_{i,k}$ corresponds to performing $\tau$ consensus iterations for the local objective values at time $k\!-\!\tau$. This element closely approximates the global objective value at time $k\!-\!\tau$ with a suitable $\tau$, because $\bm{z}^{(1)}_{i,k}(1)$ equals $\sum_{j=1}^{N}W^{\tau}_{ij}\tilde{\Phi}_{j}(u_{k-\tau}+\delta v_{k-\tau})$.
Let $\bm{Z}_{k}\in\mathbb{R}^{(\tau+1)\times N}$ be the matrix whose $i$-th column is $\bm{z}^{(1)}_{i,k}$. We denote the first row of $\bm{Z}_{k}$ by $\bm{Z}_{k}(1)\in\mathbb{R}^{N}$. For $k\geq\tau+1$, the update for all the agents can be compactly written as
$u_{k+1}=u_{k}-\eta\frac{1}{\delta}\big{(}\underbrace{\bm{Z}_{k}(1)^{\top}}_{W^{\tau}\varphi_{k\!-\!\tau}}-\underbrace{\bm{Z}_{k-1}(1)^{\top}}_{W^{\tau}\varphi_{k\!-\!\tau\!-\!1}}\big{)}\odot v_{k-\tau},$ (7)
@@
The parameter $\tau$ regulates the trade-off between the accuracy of the gradient estimate and the communication burden. On the one hand, if $\tau$ is large, then $W^{\tau}$ approximates $\frac{1}{N}\mathbf{1}\mathbf{1}^{\top}$ well, where $\mathbf{1}\in\mathbb{R}^{N}$ is an all-ones vector. Thus, the distributed update (7) owns a comparable accuracy guarantee as that of its centralized counterpart. On the other hand, a large $\tau$ indicates an increased amount of data exchanged between neighboring agents at each iteration. In practice, we can choose $\tau$ according to the bandwidth of the communication channel and the specified accuracy requirement.
@@ III-B Constrained Setting
where $\text{Proj}_{\mathcal{U}_{i}}[\cdot]$ denotes the projection onto $\mathcal{U}_{i}$, and $\Delta_{k}$ is as in (6).
@@ IV Performance Guarantees
We provide convergence guarantees when the proposed distributed model-free controller (7) is implemented in closed loop with the system (1). We use the expected distance to the unique solution $u^{*}$ to problem (4), i.e., $\mathbb{E}[\|u_{k}-u^{*}\|^{2}]$, as the convergence measure. This metric is common in the literature when the objective is strongly convex and smooth [27, 15].
@@ IV-A Unconstrained Setting
Let Assumptions 1 and 2 hold and $\mathcal{U}=\mathbb{R}^{N}$. If $0<\eta<\frac{\delta}{\sqrt{4NL_{0}^{2}\text{tr}(W^{2\tau})}}$, then for any $k>1$, $\mathbb{E}[\|\tilde{\nabla}\tilde{\Phi}(u_{k})+e_{k}\|^{2}]$, $\mathbb{E}[\|\tilde{\nabla}\tilde{\Phi}(u_{k})\|^{2}]$, and $\mathbb{E}[\|e_{k}\|^{2}]$ are bounded from above by $R$, $R_{f}$, and $R_{e}$, respectively, where $R$, $R_{f}$, and $R_{e}$ are constants give by
$\begin{split}R=&\frac{1}{\alpha}\mathbb{E}[\|\tilde{\nabla}\tilde{\Phi}(u_{0})+e_{0}\|^{2}]+16L_{0}^{2}\text{tr}[W^{2\tau}](N+4)^{2}\frac{1}{1-\alpha},\\ R_{f}=&\frac{\alpha R}{2\text{tr}[W^{2\tau}]}+8L_{0}^{2}(N+4)^{2},\\ R_{e}=&\text{tr}[(W^{\tau}-\frac{1}{N}\mathbf{1}\mathbf{1}^{\top})^{2}]R_{f},\\ \end{split}$
and the constant $\alpha=\frac{4N}{\delta^{2}}L_{0}^{2}\eta^{2}\text{tr}[W^{2\tau}]$.
@@
Next, we characterize the optimality of the closed-loop interconnection of the proposed controller (7) and system (1).
@@
The upper bound (10) suggests that the influence of the transient distance to optimality $\mathbb{E}[\|u_{\tau+1}-u^{*}\|^{2}]$ vanishes exponentially fast. The error bound in the limit (i.e., $p(\eta)/(1-\rho)$) results from the bias of gradient estimates and the consensus errors. We further show that the limiting error bound can be made arbitrarily small by tuning $\tau$.
@@
$\eta_{\epsilon}$ $=(\sqrt{(a_{2}\!+\!0.5\epsilon m)^{2}\!+\!(2a_{1}\!+\!32N^{2}L_{1}L_{0}^{2}\text{tr}(W^{2\tau}))(m\!-\!1)\epsilon}$
@@ IV-B Constrained Setting
With input constraints in place, the limiting bound on the distance to the optimal point (i.e., $\eta R^{\prime}/(1-\rho^{\prime})$) depends on the step size $\eta$ as well as the diameter of the feasible set $D_{\mathcal{U}}$.
@@ V Application in Distributed Voltage Control
When there is an unknown load change $\Delta I\in\mathbb{R}^{8}$ in $I^{*}$, we aim to control the voltage of the system (12) close to a reference value $V_{\text{m, ref}}$ while minimizing the effort associated to controllable current injection. The problem is formulated as
@@
where the input-output steady-state sensitivity matrix $H=\begin{bmatrix}I&0\end{bmatrix}\begin{bmatrix}G&-B\\ B^{\top}&-R\end{bmatrix}^{-1}\begin{bmatrix}I\\ 0\end{bmatrix}\in\mathbb{R}^{8\times 8}$, and $V_{\text{m, ref}}=HI^{*}+d$. Note that local agents do not know the dynamics model (12) or the sensitivity matrix $H$. Instead, they utilize local output measurements and collaboratively solve problem (13). In experiments, the communication network is aligned with the physical structure, and $C$, $G$, and $L$ are set as identity matrices. Further, $I^{*}$ and $\Delta I$ are all-ones vectors, and the resistance $R$ of each line is $10$. We apply Euler’s forward method to discretize the system (12) by setting the discretization step size $\epsilon=0.1$. We set the step size $\eta=0.001$ and smoothing parameter $\delta=0.002$.
@@
In the scenario where input constraints are present, the convergence results are shown in Fig. 3b. Similar to the unconstrained scenario, the distributed controller enjoys a fast convergence rate. The gap between the distributed and centralized controllers decreases as we increase $\tau$. Constraint satisfaction is illustrated by the averaged trajectory of the input $I_{c}(6)$ in Fig. 3c, where the black dashed line represents the upper bound on the input. By exchanging more data (i.e., increasing $\tau$), the distributed controller converges to the optimal point, which is at the boundary of the feasible set.

### v2

@@ I Introduction
The above OFO controllers are mostly centralized. When applied to large-scale networked systems involving interconnected subsystems, they can face issues of scalability, robustness, and privacy disclosure, which are common in large-scale centralized decision architectures [15]. First, a central processor is needed to communicate with all the subsystems and generate network-wide control inputs. Thus, there exist severe computational and communication bottlenecks for sizable networks. Second, this centralized architecture is not robust against potential failures or attacks at the central unit. Finally, since gradient-based OFO controllers (e.g., [10, 11, 1, 7]) require the system-wide input-output sensitivity to perform local iterative updates, concerns on privacy disclosure and local modeling also arise.
@@ II-A Problem Formulation
Let the steady-state input-output map of a networked system be given by a nonlinear function $h:\mathbb{R}^{N}\times\mathbb{R}^{N}\rightarrow\mathbb{R}^{N}$
@@
where $u\in\mathbb{R}^{N}$ is the input, $y\in\mathbb{R}^{N}$ is the output, $d\in\mathbb{R}^{N}$ is a constant unknown exogenous disturbance drawn from a certain distribution. The $i$-th elements of $u$, $y$, and $d$ are the local input, output, and disturbance of agent $i$, respectively, where $i=1,\ldots,N$. We focus on a per-agent scalar setting to simplify notation, however, our results readily extend to scenarios involving multiple variables.
@@
In problem (2), the global objective function $\Phi:\mathbb{R}^{N}\times\mathbb{R}^{N}\rightarrow\mathbb{R}$ is the average of all the local objectives $\Phi_{i}:\mathbb{R}\times\mathbb{R}\rightarrow\mathbb{R}$, and $u(i)$ and $y(i)$ are the input and output of agent $i$, respectively. Moreover, $\mathcal{U}_{i}\subseteq\mathbb{R}$ is the local input constraint set of agent $i$. The overall constraint set $\mathcal{U}$ is the Cartesian product of all the local sets. By incorporating (2b) into the objective $\Phi(u,y)$, problem (2) can be transformed to:
@@
where $\tilde{\Phi}(u)=\frac{1}{N}\sum_{i=1}^{N}\tilde{\Phi}_{i}(u)$, and $\tilde{\Phi}_{i}(u)\triangleq\Phi_{i}(u(i),y(i))$ is the reduced local objective. Also, $y(i)$ denotes the $i$-th element of the overall output $y=h(u,d)$. The disturbance $d$ acts as an unknown parameter of $\tilde{\Phi}(u)$. We make the following assumptions on the objectives and the network.
@@
Both Assumptions 1 and 2 are standard in the literature of network control, see [28, 29, 27] and [15, 30], respectively. For instance, the assumption that $\tilde{\Phi}_{i}$ is strongly convex can be satisfied when the system has linear dynamics and the objective $\Phi_{i}$ is a strongly convex function of $u(i)$ and $y(i)$.
@@ III Design of Distributed Model-free OFO Controllers
In contrast, we pursue an OFO controller to regulate the system (1) in a distributed manner. This controller employs a novel communication protocol, where each agent stores and exchanges a vector of historical evaluations of local objectives for tracking the global objective value. This vector is updated via average consensus in the same time scale as the local optimization iterations. In the remainder of this paper, we treat the unconstrained case of (3) (i.e., $\mathcal{U}=\mathbb{R}^{N}$) separately, to derive sharper results.
@@ III-A Unconstrained Setting
4:        $\bm{z}_{i,k}(l)=\sum_{j\in\mathcal{N}_{i}}W_{ij}\bm{z}_{j,k}(l)$, $\forall l=1,\ldots,\tau$;
5:        $\bm{z}^{(1)}_{i,k}=[\bm{z}^{\top}_{i,k},\tilde{\Phi}_{i}(u_{k}+\delta v_{k})]^{\top}$;
@@
$\Delta_{k}(i)\!=\!\begin{cases}\bm{z}^{(1)}_{i,k}(1)v^{\text{init}}_{0}(i)&\text{if $k\!=\!0$},\\ \big{(}\bm{z}^{(1)}_{i,k}(1)\!-\!\bm{z}^{(1)}_{i,k-1}(1)\big{)}v^{\text{init}}_{k}(i)&\text{if $0\!<\!k\!<\!\tau$},\\ \big{(}\bm{z}^{(1)}_{i,k}(1)\!-\!\bm{z}^{(1)}_{i,k-1}(1)\big{)}v_{k-\tau}(i)&\text{if $k\!\geq\!\tau$};\end{cases}$ (5)
7:        $\bm{z}_{i,k+1}=[\bm{z}^{(1)}_{i,k}(2),\ldots,\bm{z}^{(1)}_{i,k}(\tau+1)]^{\top}$;
@@
Let $\bm{z}_{i,k}$ be a vector containing the local objective values of agent $i$ at iteration $k$ and $\bm{z}_{i,k}(j)$ be the $j$-th entry of $\bm{z}_{i,k}$. The model-free distributed controller requires a suitable initialization of $\bm{z}_{i,0}\in\mathbb{R}^{\tau}$, where $\tau\in\mathbb{N}$ is a design parameter. For this, agent $i$ can evaluate the local objective $\tau$ times using $v^{\text{init}}_{l}(i)\sim\mathcal{N}(0,1)$ for $l=0,\ldots,\tau-1$ and stack the result in $\bm{z}_{i,0}$. Consequently, we have
@@
We summarize the detailed steps of the proposed distributed OFO controller in Algorithm 1. The shared coefficients, such as $\tau,\eta$, and $\delta$, can be either set in advance or synchronized via consensus mechanisms. At every iteration $k$, each agent $i$ generates a random exploration signal $v_{k}(i)\sim\mathcal{N}(0,1)$ synchronously and evaluates its local objective value $\tilde{\Phi}_{i}(u_{k}+\delta v_{k})=\Phi_{i}(u_{k}(i)+\delta v_{k}(i),y_{k}(i))$ with its local input $u_{k}(i)+\delta v_{k}(i)$ and local measured output $y_{k}(i)$, which captures the collective effect of the coupled dynamics. These local objective values are stored in a vector $\bm{z}_{i,k}\in\mathbb{R}^{\tau}$ of size $\tau$. Then, the agents update $\bm{z}_{i,k}$ via the average consensus step (see line 4 in Algorithm 1), where $W$ is a doubly stochastic matrix with Metropolis weights[15]. Afterward, the new local evaluation of $\tilde{\Phi}_{i}$ queues into $\bm{z}_{i,k}$. We denote this vector with $\tau+1$ elements by $\bm{z}^{(1)}_{i,k}$. The first element of $\bm{z}^{(1)}_{i,k}$ is taken out to facilitate the optimization iterations (see lines 5-7 in Algorithm 1). We illustrate the process of communicating and updating $\bm{z}_{i,k}$ in Fig. 1.
In fact, when $k\geq\tau+1$, the first element of $\bm{z}^{(1)}_{i,k}$ corresponds to performing $\tau$ consensus iterations for the local objective values at time $k\!-\!\tau$. This element closely approximates the global objective value at time $k\!-\!\tau$ with a suitable $\tau$, since $\bm{z}^{(1)}_{i,k}(1)=\sum_{j=1}^{N}W^{\tau}_{ij}\tilde{\Phi}_{j}(u_{k-\tau}+\delta v_{k-\tau})$.
We write $\bm{Z}_{k}=[\bm{z}^{(1)}_{1,k}(1),\dots,\bm{z}^{(1)}_{N,k}(1)]^{\top}$, where $\bm{Z}_{k}\in\mathbb{R}^{N}$. For $k\geq\tau+1$, the update for all the agents is
$u_{k+1}=u_{k}-\eta\frac{1}{\delta}\big{(}\underbrace{\bm{Z}_{k}}_{W^{\tau}\varphi_{k\!-\!\tau}}-\underbrace{\bm{Z}_{k-1}}_{W^{\tau}\varphi_{k\!-\!\tau\!-\!1}}\big{)}\odot v_{k-\tau},$ (6)
@@
The parameter $\tau$ regulates the trade-off between the accuracy of the gradient estimate and the communication burden. On the one hand, if $\tau$ is large, then $W^{\tau}$ approximates $\frac{1}{N}\mathbf{1}\mathbf{1}^{\top}$ well, where $\mathbf{1}\in\mathbb{R}^{N}$ is an all-ones vector. Thus, the distributed update (6) owns a comparable accuracy guarantee as that of its centralized counterpart. On the other hand, a large $\tau$ indicates an increased amount of data exchanged between neighboring agents at each iteration. In practice, we can choose $\tau$ according to the bandwidth of the communication channel and the accuracy requirement.
@@ III-B Constrained Setting
where $\text{Proj}_{\mathcal{U}_{i}}[\cdot]$ denotes the projection onto $\mathcal{U}_{i}$, and $\Delta_{k}$ is defined as in (5).
@@ IV Performance Guarantees
We provide convergence guarantees when the proposed distributed model-free controller (6) is implemented in the closed loop with system (1). We use the expected distance to the unique solution $u^{*}$ to problem (3), i.e., $\mathbb{E}[\|u_{k}-u^{*}\|^{2}]$, as the convergence measure. This metric is common in the literature when the objective is strongly convex and smooth [27, 15].
@@ IV-A Unconstrained Setting
Let Assumptions 1 and 2 hold and $\mathcal{U}=\mathbb{R}^{N}$. If $0<\eta<\frac{\delta}{\sqrt{4NL_{0}^{2}\text{tr}(W^{2\tau})}}$, then for any $k>1$, $\mathbb{E}[\|\tilde{\nabla}\tilde{\Phi}(u_{k})+e_{k}\|^{2}]$, $\mathbb{E}[\|\tilde{\nabla}\tilde{\Phi}(u_{k})\|^{2}]$, and $\mathbb{E}[\|e_{k}\|^{2}]$ are bounded from above by $R$, $R_{f}$, and $R_{e}$, respectively, where $R=\mathcal{O}(N^{2})$, $R_{f}=\mathcal{O}(N^{2})$, and $R_{e}=\mathcal{O}(N^{2},\text{tr}[(W^{\tau}-\frac{1}{N}\mathbf{1}\mathbf{1}^{\top})^{2}])$ are constants, which can be found in Appendix -B.
@@
Next, we characterize the optimality of the closed-loop interconnection of the proposed distributed controller (6) and the system (1).
@@
The inequality (9) suggests that the transient behavior of our distributed algorithm is captured by the convergence of $\mathbb{E}[\|u_{k+1}-u^{*}\|^{2}]$ with a linear rate $\rho$. Further, the asymptotic behavior reflecting the suboptimality of the limiting point is dominated by errors due to stochastic gradient estimation and consensus-based tracking mechanisms. We further show that the bound of suboptimality (i.e., $p(\eta)/(1-\rho)$) asymptotically converges to zero as $\tau$ goes to infinity.
@@
$\eta_{\epsilon}$ $=(\sqrt{(a_{2}\!+\!0.5m\epsilon)^{2}\!+\!(2a_{1}\!+\!32N^{2}L_{1}L_{0}^{2}\text{tr}(W^{2\tau}))(m\!-\!1)\epsilon}$
@@ IV-B Constrained Setting
With input constraints in place, the limiting bound on the distance to the optimal point (i.e., $\eta R^{\prime}/(1-\rho^{\prime})$) depends on the step size $\eta$ and the diameter of the feasible set $D_{\mathcal{U}}$. Compared to the unconstrained case, the distance between any two points in the feasible set is not more than $\mathcal{D_{\mathcal{U}}}$ in the constrained case, which eliminates the need for dealing with coupling transients in $u$. While the use of $\mathcal{D_{\mathcal{U}}}$ facilitates analysis, the upper bound (10) tends to be more conservative than the bound (9).
@@ V Application in Distributed Voltage Control
When there is an unknown load change $\Delta I\in\mathbb{R}^{8}$ in $I^{*}$, we aim to control the voltage of the system (11) close to a reference value $V_{\text{m, ref}}$ while minimizing the effort associated to controllable current injection. The problem is formulated as follows
@@
where the input-output steady-state sensitivity matrix $H=\begin{bmatrix}I&0\end{bmatrix}\begin{bmatrix}G&-B\\ B^{\top}&-R\end{bmatrix}^{-1}\begin{bmatrix}I\\ 0\end{bmatrix}\in\mathbb{R}^{8\times 8}$, and $V_{\text{m, ref}}=HI^{*}+d$. Note that local agents do not know the dynamics model (11) or the sensitivity matrix $H$. Instead, they utilize local output measurements and collaboratively solve problem (12). In simulations, the communication network is aligned with the physical structure, and $C$, $G$, and $L$ are set as identity matrices. Further, $I^{*}$ and $\Delta I$ are all-ones vectors, and the resistance $R$ of each line is $10$. We apply Euler’s forward method to discretize the system (11) by setting the discretization step size $\epsilon=0.1$. We set the step size $\eta=0.001$ and smoothing parameter $\delta=0.002$.
@@
For the scenario with input constraints, the convergence results are shown in Fig. 3b. Similar to the unconstrained scenario, the distributed controller enjoys a fast convergence rate. The gap between the distributed and centralized controllers decreases as we increase $\tau$. Constraint satisfaction is illustrated by the averaged trajectory of the input $I_{c}(6)$ in Fig. 3c, where the black dashed line represents the upper bound on the input. By exchanging more data (i.e., increasing $\tau$), the distributed controller converges to the optimal point, which is at the boundary of the feasible set.